P0 = R130; Ps = R 100; Pm = R200; Q1 = 1000; Q0 = 1200

Q.9.3.1 Calculate supplier’s revenue before tax. (2)

Q.9.3.2 Calculate the supplier’s revenue after tax. (2)

Q.9.3.3 Calculate the total tax revenue collected by the government. (2)

Q.9.3.4 Calculate the total tax paid by the consumers. (2)

Q.9.3.5 Calculate the total tax paid by the suppliers. (2)

Q.9.3.6 Compare the results in 9.3.4 and 9.3.5 above and give a logical

explanation for the variance.

(4)


1
Expert's answer
2021-08-25T18:05:03-0400

Q.9.3.1

Before tax :-

Price = P0 = R130

Quantity = Q0 = 1200

Hence, before tax revenue for the supplier

"= 130 \u00d7 1200 = R156,000"


Q.9.3.2

After tax :-

Price for supplier = PS = R100

Quantity = Q1 = 1000

After tax revenue "= 100 \u00d7 1000 = R100,000"


Q.9.3.3

Value of tax "=P_m - P_s = 200 - 100 = R100"

After tax quantity = 1000

Hence, total tax revenue collected by the government "= 100 \u00d7 1000 = R100,000"


Q.9.3.4

Total tax paid by the consumers

"= (P_m - P_0) × Q_1\\

= (200 - 130) × 1000\\

= R 70,000"

 

Q.9.3.5

Total tax paid by the suppliers

"= (P_0 - P_S) × Q_1\\

= (130 - 100) × 1000\\

= R 30,000"


Q.9.3.6

consumers pay more tax because demand is inelastic than supply.